OpenStudy (anonymous):

which number classified as a rational number? square root17 square root23 square root9 square root 2

OpenStudy (anonymous):

\[\sqrt{9} = ??\]

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Rational Numbers are the numbers that can be represented in the form of \(\frac{p}{q}\), where \(p\) and \(q\) are integers..

OpenStudy (anonymous):

so is 9 right?

OpenStudy (anonymous):

So, \[\sqrt{9} = 3 = \frac{3}{1} = \frac{p}{q}\]

OpenStudy (anonymous):

9 is not right, but \(\sqrt{9}\) is.. :P

OpenStudy (anonymous):

please help me